HST Status Mar 09, 2011

3 views
Skip to first unread message

Dr. Paul Scowen

unread,
Mar 23, 2011, 12:28:35 PM3/23/11
to

Earlier reports Convert Day of Year

The Hubble Space Telescope

Unlocking the Secrets of the Universe

Daily Report #5351

Period Covered:
07:00 pm March 8, 2011 - 06:59 pm March 9, 2011
(DOY 68/0000z - 68/2359z)

Flight Operations Summary

Significant Spacecraft Anomalies: (The following are preliminary reports
of potential non-nominal performance that will be investigated.)

HSTARs:
12757 - SCI: Engineering data was lost from D067 17:18:58z to 17:42:19z (23 Minutes 21 seconds).
There was astrometry ongoing during this time frame.@068/1718z

COMPLETED OPS REQUEST:
19050-1 - Dump COS EEPROM (4.12A)@68/1918z

COMPLETED OPS NOTES:
none

FGS ACQ STATUS:
		    SCHEDULED       SUCCESSFUL
FGS GSAcq		0		0
FGS REAcq		0		0
OBAD with Maneuver	0		0

SIGNIFICANT EVENTS:
The COS FUV HV and OSM G130M (SCR 339 and 342) FSW patches has been confirmed via a COS EEPROM 
memory dump. The SI FSW team has analyzed the dump and confirmed no differences when compared to 
the LV0044 ground reference image.   
  
FLASH COS FUV HV FSW patches successful At ~ 066/00:15 the COS FUV segment B nominal operating HV 
preset was successfully updated in Executive RAM and copied to EEPROM. Additionally the maximum FUV 
HV FSW limits which were already in Exec RAM were copied to EEPROM. Segment B nominal operating 
voltage was increased from 167 to 175 counts (delta ~125.6 volts). These changes were monitored via 
COS memory monitors. These updates are documented in COS SCR 339. The updated segment B nominal 
operating voltage preset will not be used by the FUV until it is loaded to the DCE via the LDCECONF 
macro which occurs 11.067 19:31:06.000. The new G130M central wavelength positions were also 
updated in Exec RAM and copied to EEPROM.



Program Principal Investigator Program Title
Science Observations Scheduled
12066 Marc Postman, Space Telescope Science Institute Through a Lens, Darkly - New Constraints on the Fundamental Components of the Cosmos
12068 Marc Postman, Space Telescope Science Institute Through a Lens, Darkly - New Constraints on the Fundamental Components of the Cosmos
12166 Harald Ebeling, University of Hawaii A Snapshot Survey of The Most Massive Clusters of Galaxies
12212 D. Crenshaw, Georgia State University Research Foundation What are the Locations and Kinematics of Mass Outflows in AGN?
12299 Michael Eracleous, The Pennsylvania State University Spectroscopic Signatures of Binary and Recoiling Black Holes
Calibration Observations Scheduled
12334 Jasonjot Kalirai, Space Telescope Science Institute UVIS & IR Photometric Zero Points
12342 Tiffany Borders, Space Telescope Science Institute WFC3 UVIS CCD Daily Monitor
12379 Kai Noeske, Smithsonian Institution Astrophysical Observatory UVIS CTE Monitor: Star Clusters
12380 Bryan Hilbert, Space Telescope Science Institute Guard Darks
12397 David Golimowski, Space Telescope Science Institute CCD Daily Monitor {Part 2}
12400 Michael Wolfe, Space Telescope Science Institute Dark Monitor Part 1
12402 Michael Wolfe, Space Telescope Science Institute CCD Bias and Read Noise Monitor-Part 1

Reply all
Reply to author
Forward
0 new messages